An alien from another system has a weight of 1024 N when he is on his home planet where g = 16.5 m/s2. What is his weight on Earth? 


1
Expert's answer
2021-08-10T08:01:57-0400

The weight

"W=mg"

Hence,

"W_{\rm Earth}=W_P\times \frac{g_{\rm Earth}}{g_P}\\
=1024\:\rm N\times \frac{9.8\:\rm m/s^2}{16.5\:\rm m/s^2}=608\:\rm N"